Lulabalula:

Maria just got a raise at work. She was making $9 per hour, but now she makes 150% of that amount. Explain how you would find the amount of money Maria now makes per hour.

Lulabalula:

Using Multiplication to Find Percents

Lulabalula:

@carmelle @supie @MrMudd183

Lulabalula:

Should i just close this?

Ari1234442:

NO!!!!!!

Lulabalula:

ok

carmelle:

150% = 1.5 So, we have to do 9 x 1.5 When we multiply, what do we get?

carmelle:

@Lulabalula

Ari1234442:

I found out the answer

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Ari1234442 I found out the answer \(\color{#0cbb34}{\text{End of Quote}}\) dm me the answer u got

Lulabalula:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle @Lulabalula \(\color{#0cbb34}{\text{End of Quote}}\) sorry wasn't paying attention was trying to figure it out

Ari1234442:

What do you think the answer is

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lulabalula \(\color{#0cbb34}{\text{Originally Posted by}}\) @carmelle @Lulabalula \(\color{#0cbb34}{\text{End of Quote}}\) sorry wasn't paying attention was trying to figure it out \(\color{#0cbb34}{\text{End of Quote}}\) its ok. Now, we 9 x 1.5

Lulabalula:

13.5

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lulabalula 13.5 \(\color{#0cbb34}{\text{End of Quote}}\) Based on what I did, I would say u are correct, but lets see what @Ari1234442 says

Lulabalula:

ok

carmelle:

ok

Ari1234442:

I think it is 22.5. Let me show you

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Ari1234442 I think it is 22.5. Let me show you \(\color{#0cbb34}{\text{End of Quote}}\) ok

Lulabalula:

are you going to say it?

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lulabalula are you going to say it? \(\color{#0cbb34}{\text{End of Quote}}\) idk...

Ari1234442:

9/x=100/250 then 9.250/100=x=22.5

carmelle:

9.250/100 = 0.0925 :/

Ari1234442:

it is 9.250 over 100

carmelle:

ik...thats what I did

carmelle:

Honestly, its highly unlikely for someone to go from working $9 an hour to $22.... hypothetically speaking...

carmelle:

I feel like 13.5 would fit better.. but its just my opinion

Lulabalula:

13.5 was right

Ari1234442:

OK sorry

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lulabalula 13.5 was right \(\color{#0cbb34}{\text{End of Quote}}\) ok

Ari1234442:

Math is not my thing

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Ari1234442 Math is not my thing \(\color{#0cbb34}{\text{End of Quote}}\) lol its ok (:

Lulabalula:

same

Ari1234442:

Anything else

Lulabalula:

i opened a new one

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lulabalula i opened a new one \(\color{#0cbb34}{\text{End of Quote}}\) ok I'll try to help

Ari1234442:

Where is it